A debt of $2,000 will be repaid by monthly payments of $600 for as long as necessary, the first payment to be made at the end of 6 months. If interest is at j12 = 9%, find the size of the debt at the end of 5 months and make out the complete amortization schedule starting at that time.
Answer #1

Annual interest rate=9%

Calculation of monthly interest rate:

Let monthly interest rate be x.

so,

(1+x)^12=1.09

1+x=1.09^(1/12)

x=.007207 or .7207% per month.

Size of debt at the end of 5 month=2000*(1.007207)^5

=2073.12

Amortisation schedule:

Rate of interest 0.007207
Month end Loan Outstanding Interest Installment Amount Principal Repaid Loan Balance
6 2073.12 14.94098 600 585.0590242 1488.060976
7 1488.060976 10.72446 600 589.2755445 898.7854313
8 898.7854313 6.477547 600 593.5224534 305.2629779
9 305.2629779 2.20003 600 597.7999697 -292.5369918

at the end of 9th month loan becomes negative hence repayment date should be somthing in between 8th and 9th month end. which can be found by using nper function in excel as follows:

Function Arguments NPER Rate 0.007207 Pmt 600 Pv -2073.12 B 6 6 6 = 0.007207 = 600 = -2073.12 = number Fv Type 0 3.511541207

answer is 3.51 periods means loan repayment shall be done completely after 9.51 months (i.e. 6+3.51)

=